OpenStudy (anonymous):

Medal for the first person who helps most actual help! 4.Add and answer in simplest fractional form -4/3 + 7/8= A.-11/24 B.3/5 C.3/24 5.8/9 - 12/9 = A.-4/9 B.4/9 C.-4/18 6.-1/12 - 3/4 = A.-1/3 B.-5/6 C.2/3 7.Convert each subtraction problem to addition and type in the matching number. A.-3/4 + 2/3 B.-3/4 + -2/3 C.3/4 + -2/3 D.3/4 + 2/3 *3/4 - 2/3 = *3/4 - -2/3 = *-3/4 - 2/3 = *-3/4 - -2/3 = 8.Add and answer in simplest fractional form. 3/8 + -1/8 A.1/8 B.2/16 C.1/4 9.Add and answer in simplest fractional form. -1/5 + -3/4 = A.-4/9 B.-19/20 C.-4/20 10.7/8 - (-3/16) = A.11/16 B.9/8 C.17/16

jhonyy9 (jhonyy9):

4. A

jhonyy9 (jhonyy9):

5. A

jhonyy9 (jhonyy9):

6. B

OpenStudy (anonymous):

\[-4/3 + 7/8=(-32/24)+(21/24)=-11/24\]

jhonyy9 (jhonyy9):

8. C

OpenStudy (anonymous):

\[-1/8+3/8=2/8=1/4\]

jhonyy9 (jhonyy9):

9. B

jhonyy9 (jhonyy9):

10. C

OpenStudy (anonymous):

\[-1/5 + -3/4 =(-4/20)+(-15/20)=-19/20\]

OpenStudy (anonymous):

\[8/9 - 12/9 =-4/9\]

OpenStudy (anonymous):

\[-1/12 - 3/4 =(-1/12)-(-9/12)=-10/12=-5/6\]